### v1

# Task-Unaware Lifelong Robot Learning with Retrieval-based Weighted Local Adaptation

Abstract
Real-world environments require robots to continuously acquire new skills while retaining previously learned abilities, all without the need for clearly defined task boundaries. Storing all past data to prevent forgetting is impractical due to storage and privacy concerns. To address this, we propose a method that efficiently restores a robot’s proficiency in previously learned tasks over its lifespan. Using an Episodic Memory (EM), our approach enables experience replay during training and retrieval during testing for local fine-tuning, allowing rapid adaptation to previously encountered problems without explicit task identifiers. Additionally, we introduce a selective weighting mechanism that emphasizes the most challenging segments of retrieved demonstrations, focusing local adaptation where it is most needed. This framework offers a scalable solution for lifelong learning in dynamic, task-unaware environments, combining retrieval-based adaptation with selective weighting to enhance robot performance in open-ended scenarios.
1 Introduction
Figure 1: Method Overview. Our approach addresses the challenge of lifelong learning without distinct task boundaries. To emulate human learning patterns, we propose a method consisting of three phases: learning, reviewing, and testing. In the learning phase, the robot is exposed to various demonstrations, storing a subset of this data as episodic memory $\mathcal{M}$. During the reviewing phase, the method retrieves the most relevant data to fine-tune the policy network, enhancing performance in the final testing phase.
Lifelong learning seeks to endow neural networks with the ability to continually acquire new skills while retaining previously learned knowledge. This balance between stability and plasticity is crucial as models face sequences of tasks over time. While significant progress has been made in applying lifelong learning to domains such as computer vision (Huang et al., 2024; Du et al., 2024) and natural language processing (Shi et al., 2024; Razdaibiedina et al., 2023), the challenges are more pronounced in robotics. Robots are expected to adaptively learn and solve unseen tasks throughout their operational lifespan (Thrun & Mitchell, 1995). Their interactions with dynamic environments introduce complexities absent in static data domains; a single misstep in task execution can result in complete failure. Moreover, robotics is constrained by limited data availability due to the expense and complexity of real-world interactions (Zhu et al., 2022; Du et al., 2023). These factors not only intensify the difficulty of continual learning in robotics but also demand more robust lifelong learning capabilities.
Existing methods for lifelong robot learning typically require robots to learn a sequence of tasks, each distinguished by domain, scenario, scene, or task goals (Liu et al., 2024; Yang et al., 2022; Wan et al., 2024; Parakh et al., 2024). In those settings, robots often depend on specific task identifications with clear boundaries — usually provided as task IDs or explicit descriptions — to specify the task they are working on (Liu et al., 2023). However, in dynamic real-world settings, it is impractical to predefine tasks or assign specific IDs, as robots are likely to encounter a vast array of unpredictable situations, with tasks that may be subdivided into smaller components of varying granularity. Therefore, approaches that rely on specific task identifications with clear boundaries are unrealistic and unscalable (Koh et al., 2021).
To address these challenges, we propose a novel task-unaware lifelong robot learning framework with visuomotor policies, utilizing vision perceptions as well as diverse paraphrased language descriptions. This framework enables robots to continually learn and adapt without explicit task IDs. We employ our method in manipulation scenarios based on the LIBERO benchmark (Liu et al., 2024). Our approach leverages pre-trained models to generate consistent embeddings across different tasks and training phases, thereby mitigating the embedding drift that often occurs in sequential learning scenarios (Liu et al., 2023; Kawaharazuka et al., 2024). We adopt Experience Replay (ER) baseline (de Masson D’Autume et al., 2019) to rehearse samples from previous tasks, helping to maintain learned skills and reduce forgetting.
Despite these measures, some degree of forgetting remains inevitable due to the multitasking nature of lifelong learning and the robot’s limited access to previous demonstrations. Drawing inspiration from human learning processes — where individuals revisit tasks they once knew but have forgotten details — we introduce an efficient local adaptation mechanism. Humans often perform quick reviews using limited resources and try to retrieve memory to rebuild their knowledge, allowing them to efficiently regain proficiency without relearning all aspects of the task (Sara, 2000). Similarly, our mechanism enables the robot to locally adapt to previously encountered problems rapidly and regain skills through fast fine-tuning, using the same episodic memory employed for experience replay during training.
Given the indistinct task boundaries, we leverage retrieval-based mechanisms (Du et al., 2023; van Dijk et al., 2024; de Masson D’Autume et al., 2019) to retrieve data most similar to the current task based on vision and language input similarities. To adapt the model effectively — especially focusing on the most challenging phases where the robot’s performance deviates — we first perform a few episodes of rollouts to obtain ”feedback” on the model’s performance before local adaptation: these rollouts are then used for automatic selective weighting by comparing them with the retrieved demonstrations without human intervention (Spencer et al., 2022; Mandlekar et al., 2020). The weighted samples facilitate the local adaptation phase, thereby improving performance.
In summary, the key contributions of our solution are:
  • Retrieval-Based Adaptation for Blurred Task Boundaries: During testing, relevant past demonstrations are retrieved from episodic memory to adapt the neural network locally, enabling the robot to quickly regain proficiency on previously encountered tasks without relying on explicit task boundaries.
  • Selective Weighting Mechanism: A weighting mechanism emphasizes the most challenging segments of the retrieved demonstrations, optimizing real-time adaptation.
  • Paradigm for Memory-Based Lifelong Robot Learning: We demonstrate that our approach can be applied to different memory-based robotic lifelong learning algorithms during test time, serving as a paradigm for skill restoration.
This framework allows robots to continually learn and adapt in dynamic environments without requiring predefined task identifiers or boundaries, making it highly practical and scalable for real-world applications. By combining retrieval-based adaptation with selective weighting, our method offers a robust solution to the challenges of lifelong robot learning in open-ended settings.
2 Related Work
2.1 Lifelong Robot Learning
A key challenge in lifelong robot learning is catastrophic forgetting, where learning new tasks adversely affects performance on previously learned tasks (Parisi et al., 2019). Traditional lifelong learning methods often rely on explicit task identifiers or clear task boundaries to structure the learning process (Wan et al., 2024; Xie & Finn, 2022), such as Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017) and PackNet (Mallya & Lazebnik, 2018). In real-world robotic applications, robots operate in dynamic environments where tasks are not clearly segmented, making explicit task identifiers impractical (Kim et al., 2024).
Recent efforts in lifelong reinforcement learning (Xie & Finn, 2022) and areas like task and motion planning (Mendez-Mendez et al., 2023), online object model reconstruction (Lu et al., 2022), interactive instruction-following agents (Kim et al., 2024), multi-task learning (Wang et al., 2022), interactive imitation learning (Spahn et al., 2024), and SLAM (Yin et al., 2023; Gao et al., 2022; Vödisch et al., 2022) show progress. Robot manipulation skills also evolve through interactions, aiding adaption when task executions fail (Parakh et al., 2024). Memory-based algorithms and selective weighting (Sun et al., 2022; Koh et al., 2021; Shim et al., 2021; Aljundi et al., 2019) enhance learning by prioritizing informative samples. Replay buffer methods (He et al., 2020; Mai et al., 2021; Caccia et al., 2021) have demonstrated success. However, there is still a lack of progress in settings where the model is unaware of task boundaries during both training and inference (Lee et al., 2020; Chen et al., 2020; Ardywibowo et al., 2022).
A benchmark for lifelong robot learning, particularly focusing on manipulation tasks, has been introduced in LIBERO (Liu et al., 2024). Methods such as TAIL (Liu et al., 2023) rely on specific task identifiers, which can be limiting in dynamic environments, while Lotus (Wan et al., 2024) involves a pretraining phase to establish an initial skill set, providing a foundation for further continuous learning.
2.2 Task-unaware Continual Learning
Despite the success of continual learning with clearly labeled task sequences, there still remains a gap in progress within settings where the model is unaware of task boundaries both in training and inference, an online situation more reflective of real-world scenarios. Many attempts (Lee et al., 2020; Chen et al., 2020; Ardywibowo et al., 2022) focus on learning specialized parameters using expanding network structures. Memory-based algorithms remain effective by prioritizing informative samples (Sun et al., 2022), removing less important training samples (Koh et al., 2021), improving decision boundaries (Shim et al., 2021), and increasing gradient diversity (Aljundi et al., 2019). Methods aiming to exploit replay buffer (He et al., 2020; Mai et al., 2021; Caccia et al., 2021) have also demonstrated notable success.
2.3 Information Retrieval for Robotics
Information retrieval techniques have been used to optimize robotic behaviors by retrieving relevant actions from memory in novel tasks (Du et al., 2023). For example, path following based on image retrieval improves visual navigation (van Dijk et al., 2024), and incremental learning helps humanoid robots adapt to new environments by recalling past behaviors (Bärmann et al., 2023). Retrieval has also enabled skill transfer from videos (Papagiannis et al., 2024) and affordance transfer for zero-shot manipulation (Kuang et al., 2024), allowing robots to manipulate objects without prior training.
2.4 Robot Learning with Adaptation
Robots have learned to adapt to dynamic environments, such as agile flight in strong winds (O’Connell et al., 2022) and quadruped locomotion adaptation through test-time search (Peng et al., 2020). Meta-learning allows fast adaptation to new tasks (Kaushik et al., 2020; Nagabandi et al., 2018), while efficient adaptation techniques enable robots to generalize from limited data (Julian et al., 2020). Despite the success of retrieval-based adaptation and selective weighting, combining these methods for lifelong learning in open-ended environments remains an open challenge.
Figure 2: Policy Backbone Architecture used in Training and Testing. We input various data modalities into the system, including demonstration images, language descriptions, and the robot arm’s proprioceptive input (joint and gripper states). Pretrained R3M (Nair et al., 2022) and SentenceSimilarity (2024) models process the image and language data respectively. Along with the proprioceptive states processed by an MLP, the embeddings are concatenated and passed through a Transformer to generate temporal embeddings. A GMM (Gaussian Mixture Model) is then used as the policy head to sample actions for the robot. Throughout both training and testing, we utilize episodic memory to store a subset of demonstrations gathered throughout the training process.
3 Preliminary
Our robot utilizes a visuomotor policy learned through behavior cloning to execute manipulation tasks by mapping sensory inputs and task descriptions to motor actions. In a task-unaware lifelong learning setting, we adopt a continual learning framework where task boundaries are blurred by employing multiple paraphrased descriptions to define task goals, rather than relying on explicit task identifiers. This approach enhances the policy’s ability to generalize across varied instructions and tasks.
The policy is trained by minimizing the discrepancy between the predicted actions and the expert actions derived from demonstrations. Specifically, we optimize the following loss function across a sequence of tasks $\{\mathcal{T}_{k}\}$ with corresponding demonstrations $\mathcal{D}_{k}=\{\tau^{k}_{1},\ldots,\tau^{k}_{N}\}$. Notably, $\mathcal{D}_{k}$ is not fully accessible for $k<K$ due to the use of experience replay data from Episodic Memory $\mathcal{M}$:
$\theta^{*}=\arg\min_{\theta}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(o_{t},a_{t})\sim\mathcal{D}_{k},\,g\sim G_{k}}\left[\sum_{t=0}^{l_{k}}\mathcal{L}\left(\pi_{\theta}(o_{\leq t},g),\,a_{t}\right)\right],$ (1)
where $\theta$ denotes the model parameters, $l_{k}$ represents the number of samples for task $k$, $o_{\leq t}$ denotes the sequence of observations up to time $t$ in demonstration $n$ (i.e., $o_{\leq t}=(o_{0},o_{1},\ldots,o_{t})$), and $a_{t}$ is the expert action at time $t$. The set $G_{k}$ comprises paraphrased goal descriptions for task $\mathcal{T}_{k}$, with $g$ being a sampled goal description from $G_{k}$. The policy output, $\pi_{\theta}(o_{\leq t},g)$, is conditioned on both the observation sequence and the goal description.
By optimizing this objective function, the policy effectively continues learning new tasks and skills in its life span, without the need for explicit task labels, thereby facilitating robust and adaptable task-unaware continual learning.
4 Retrieval-based Weighted Local Adaptation for Lifelong Robot Learning
In this section, we outline our proposed method depicted in Figure 1. To effectively interact with complex physical environments, the network integrates multiple input modalities, including visual inputs from workspace and wrist cameras, proprioceptive inputs of joint and gripper states, and paraphrased task descriptions.
Instead of training all modules jointly in an end-to-end manner, we employ pretrained visual and language encoders that leverage prior semantic knowledge. Pretrained encoders enhance performance on downstream manipulation tasks (Liu et al., 2023) and are well-suited to differentiate between various scenarios and tasks without relying on explicit task identifiers or clear task boundaries. Their consistent representations when new tasks continue to come is essential for managing multitask problems and retrieving relevant data to support our proposed local adaptation during test time.
When learning new tasks, the robot preserves previously acquired skills by replaying prior manipulation demonstrations stored in an episodic memory $\mathcal{M}$, which contains a small subset of previous task demonstrations (Chaudhry et al., 2019). Trained with the combined data from the latest scenarios and episodic memory $\mathcal{M}$, the model can acquire new skills while mitigating catastrophic forgetting of old tasks, thereby maintaining a balance between stability and plasticity (Wang et al., 2024). Figure 2 illustrates the network architecture, and implementation details are provided in Section A.2.
4.1 Data Retrieval
During deployment, we first retrieve the most relevant demonstrations from episodic memory $\mathcal{M}$ based on similarity to the current scenario. Due to the blurred task boundaries, some tasks share similar visual observations but differ in their task objectives, while others have similar goals but involve different backgrounds, objects, etc. To account for these variations, we compare both visual inputs from the workspace camera (Du et al., 2023) and task descriptions (de Masson D’Autume et al., 2019) using $L_{2}$ distances of their embeddings. The retrieval process follows a simple rule:
$\mathcal{D}_{R}=\alpha_{v}\cdot\mathcal{D}_{v}+\alpha_{l}\cdot\mathcal{D}_{l},$ (2)
where $\mathcal{D}_{R}$ is the weighted retrieval distance, $\mathcal{D}_{v}$ represents the distance between the embeddings of the scene observation from the workspace camera, and $\mathcal{D}_{l}$ depicts the distance between the task description embeddings. The parameters $\alpha_{v}$ and $\alpha_{l}$ control the relative importance of visual and language-based distances. Based on the distances $\mathcal{D}_{R}$, the most relevant demonstrations can be retrieved from $\mathcal{M}$.
4.2 Weighted Local Adaptation
Figure 3: Trajectory and Weighting Visualizations. To identify the point of failure, we compute the similarity between the retrieved demonstrations and failed trajectories at each frame. Once the separation segment is detected, higher weights are assigned to the frames in the segment of retrieved demonstrations during local adaptation.
4.2.1 Learn from Errors by Selective Weighting
To make the best use of the limited data, we enhance their utility by assigning weights to critical or vulnerable segments in each retrieved demonstration. Specifically, before testing, the robot performs several rollouts on the encountered task using the existing model trained during the lifelong learning phase. This procedure allows us to evaluate the model’s performance and identify any forgetting effects, akin to a preliminary quiz before the final exam (as illustrated in step 2, the reviewing phase in Figure 1).
When failed trajectories are identified, we compare each image in the retrieved demonstrations against all images from the failed trajectories using the $L_{2}$ distances of their embeddings. This comparison yields a distance vector for each demonstration, where each value represents the minimal distance between a demonstration frame and all images from the failed rollouts. This metric determines whether a particular frame has occurred during the rollout. Through this process, we identify the Separation Segment — frames in the demonstrations where the behavior deviates from what was executed during the failed rollouts (see Figure 3). Since these Separation Segments highlight behaviors that should have occurred but did not, we consider them vulnerable segments that contribute to the failure. We assign higher weights to these frames which will scale the losses during local adaptation. Detailed heuristics and implementation specifics are provided in Appendix A.4.
4.2.2 Adaptation with Fast Finetuning
Finally, we fine-tune the network’s parameters to better adapt to the current task using the retrieved demonstrations, focusing more on the difficult steps identified through selective weighting. Notably, the episodic memory $\mathcal{M}$ contains the same data used during training for experience replay and during deployment for local adaptation. No additional demonstrations are available to the robot at test time. Despite this limited data, our experiments demonstrate that the model can effectively recover learned skills and improve its performance across various tasks. Overall, the proposed weighted adaptation is formalized as follows:
$\mathbf{\theta}^{*}=\arg\min_{\mathbf{\theta}}\sum_{n=1}^{\tilde{N}}\sum_{t=1}^{l_{n}}w_{t,n}\mathcal{L}\left(\pi_{\theta}(o_{\leq t,n},g_{n}),a_{t,n}\right)$ (3)
where $\tilde{N}$ is the number of retrieved demonstrations, $l_{n}$ is the length of demonstration $n$, and $w_{t,n}$ is the weight assigned to sample $t$ in demonstration $n$. The variables $o_{\leq t,n}$ and $a_{t,n}$ denote the sequence of observations up to time $t$ and the corresponding expert action, respectively, while $g_{n}$ is the goal description for demonstration $n$. The parameter $\mathbf{\theta}$ represents the network’s parameters before adaptation.
5 Experiments
We conduct a comprehensive set of experiments to evaluate the effectiveness of our proposed retrieval-based weighted local adaptation method for lifelong robot learning. Specifically, our experiments aim to address the following key questions:
  1. 1.
    Effect of Blurry Task Boundaries: How do blurry task boundaries influence the model’s performance and data retrieval during testing?
  2. 2.
    Advantages of Retrieval-Based Adaptation: Does retrieval-based weighted local adaptation enhance the robot’s performance across diverse tasks?
  3. 3.
    Impact of Selective Weighting: Is selective weighting based on rollout errors effective in improving task performance?
  4. 4.
    Generalizability: Can our method be applied to different memory-based lifelong robot learning approaches, serving as a paradigm that enhances the performance during test time by restoring previous knowledge and skills?
  5. 5.
    Robustness: Due to blurry task boundaries and retrieval imprecision, the retrieved demonstrations may not necessarily belong to the same task. How resilient is our method to inaccuracies in memory retrieval?
5.1 Experimental Setup
5.1.1 Benchmarks
We evaluate our proposed methods using LIBERO (Liu et al., 2024): libero_spatial, libero_object, libero_goal, and libero_different_scenes. These environments feature a variety of objects and layouts. The first three benchmarks all include $10$ distinct tasks, each with up to $50$ demonstrations collected in simulation with different initial states of objects and the robot. Specifically, libero_different_scenes is created from LIBERO’s provided LIBERO_90, which encompasses $20$ tasks from distinct scenes.
For each task, we paraphrased the assigned single goal description into diverse descriptions to obscure task boundaries. These enriched descriptions were generated by rephrasing the original task descriptions from the benchmark using a large language model provided by Phi-3-mini-4k-instruct Model (mini-4k instruct, 2024), ensuring consistent meanings while varying phraseology and syntax. Please see Section A.3 for more details.
5.1.2 Baselines
We evaluate our proposed method against the following baseline approaches:
  1. 1.
    Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017): A regularization-based approach that constrains updates to the network’s parameters to prevent catastrophic forgetting of previously learned tasks.
  2. 2.
    Experience Replay (ER) (Chaudhry et al., 2019): A core component of our training setup, ER utilizes stored episodic memory to replay past demonstrations, helping the model maintain previously acquired skills and mitigate forgetting. As a baseline, we evaluate the standalone performance of ER without additional retrieval-based adaptation techniques.
  3. 3.
    Average Gradient Episodic Memory (AGEM) (Hu et al., 2020): Employs a memory buffer to constrain gradients during the training of new tasks, ensuring that updates do not interfere with performance on earlier tasks.
  4. 4.
    AGEM with Weighted Local Adaptation (AGEM-WLA): An extension of AGEM that incorporates weighted local adaptation during the testing phase, enhancing the model’s ability to adapt to specific tasks based on retrieved demonstrations. This allows us to assess the generalizability of our proposed method as a paradigm framework on other memory-based lifelong learning approaches.
  5. 5.
    PackNet (Mallya & Lazebnik, 2018): An architecture-based lifelong learning algorithm that iteratively prunes the network after training each task, preserving essential nodes while removing less critical connections to accommodate subsequent tasks. However, its pruning and post-training phases rely heavily on clearly defined task boundaries, making PackNet a reference baseline when task boundaries are well-defined.
5.1.3 Metrics
Our primary focus is on the success rate of task execution, as it is a crucial metric for manipulation tasks in interactive robotics. Consequently, we adopt the Average Success Rate (ASR) as our primary evaluation metric to address the challenge of catastrophic forgetting within the lifelong learning framework, evaluating success rates on three random seeds across all diverse tasks within the same benchmark.
5.1.4 Model, Training, and Adaptation
As illustrated in Figure. 2, our model utilizes pretrained encoders for visual and language inputs: R3M (Nair et al., 2022) for visual encoding, Sentence Similarity model (SS Model) (SentenceSimilarity, 2024) for language embeddings, and a trainable MLP-based network to encode proprioceptive inputs. Embeddings from ten consecutive time steps are processed through a transformer-based temporal encoder, with the resulting output passed to a GMM-based policy head for action sampling. Specifically, R3M, a ResNet-based model trained on egocentric videos using contrastive learning, captures temporal dynamics and semantic features from scenes, while Sentence Similarity Model captures semantic relationships in task descriptions, enabling the model to differentiate between various natural language instructions.
We trained the model sequentially on multiple tasks, with each task trained for $50$ epochs. For later tasks, we used experience replay, storing $8$ demonstrations per task from the episodic memory $\mathcal{M}$. Every 10 epochs, we evaluated the model on the current task over $20$ episodes, and the model with the highest ASR would be saved.
During testing, the agent performs 10 rollout episodes (referred to as a quiz phase) without adaptation, to assess initial performance and potential forgetting. We then retrieve the top $10\%$ most similar demonstrations from $\mathcal{M}$ using visual and language embeddings. Then, the model is fine-tuned for $20$ epochs with those weighted retrieved demonstrations. Finally, the adapted model is evaluated over $20$ episodes again (the final testing phase) to assess performance improvements. For experiments on all benchmarks, we train and test the models with three random seeds ($1$, $21$, and $42$) to reduce the impact of randomness.
5.2 Results
5.2.1 Comparison with Baselines
To address Question 2, we compared our proposed method, Weighted Local Adaptation (ER-WLA), with all baseline approaches. As shown in Table 1, ER-WLA consistently outperforms baselines of EWC, AGEM, ER, and AGEM-WLA, which do not rely on clear task boundaries. By incorporating local adaptation during test time — our method mirrors how humans review and reinforce knowledge when it is partially forgotten — the continually learning robot could also regain its proficiency on previous tasks.
In contrast, PackNet serves as a reference method, as it requires well-defined task boundaries. However, as the number of tasks increases, the network’s trainable capacity under PackNet diminishes, leaving less flexibility for future tasks. This limitation becomes evident in the libero_different_scenes benchmark, which includes 20 tasks. PackNet’s success rate drops significantly for later tasks, resulting in poor overall performance and highlighting its constraints on plasticity compared with our proposed ER-WLA approach.
Additionally, when we applied WLA to the AGEM baseline (resulting in AGEM-WLA), it also improved its performance, demonstrating the effectiveness of our method as a paradigm for memory-based lifelong robot learning methods. These findings also support our conclusions regarding Question 4.
Table 1: Comparison with Baselines, the average success rates and standard deviations across various baselines are shown below. We provide PackNet’s performance on the right as a reference point for cases where task boundaries are accessible. Both EWC and vanilla AGEM demonstrate weak performance across all benchmarks, while ER performs better due to memory replay. Under our weighted local adaptation (WLA) paradigm, the WLA-enhanced versions of ER and AGEM show significant improvements over their vanilla counterparts, highlighting the effectiveness of WLA.
Method EWC AGEM AGEM-WLA ER ER-WLA PackNet
libero_spatial 0.0 $\pm$ 0.0 7.33 $\pm$ 14.25 35.83 $\pm$ 15.71 15.67 $\pm$ 13.50 39.83 $\pm$ 19.85 53.17 $\pm$ 24.72
libero_object 1.50 $\pm$ 3.26 27.17 $\pm$ 22.77 51.17 $\pm$ 24.13 56.50 $\pm$ 19.88 62.33 $\pm$ 18.69 73.77 $\pm$ 16.97
libero_goal 0.33 $\pm$ 1.83 10.83 $\pm$ 16.03 58.67 $\pm$ 25.93 52.33 $\pm$ 22.16 62.33 $\pm$ 28.75 66.33 $\pm$ 24.88
libero_different_scenes 2.58 $\pm$ 8.98 20.43 $\pm$ 25.55 41.75 $\pm$ 33.50 34.08 $\pm$ 28.55 45.17 $\pm$ 31.86 32.92 $\pm$ 44.03
5.2.2 Ablation Studies
We performed two ablation studies to validate the effectiveness of our implementation choices and address Questions 1, 3, and 5.
Selective Weighting.
In the first ablation, we evaluated the impact of selective weighting on libero_spatial, libero_object, and libero_goal benchmarks to demonstrate its importance for effective local adaptation. We compared two variants of our method: 1) ER-ULA, which applies uniform local adaptation without selective weighting, adapting retrieved demonstrations uniformly; 2) ER-WLA, which incorporates selective weighting during test-time adaptation. Both methods are trained with experience replay.
Since early stopping during local adaptation at test time is infeasible, and training can be unstable, particularly regarding manipulation success rates, we conducted adaptation using three different numbers of epochs — $15$, $20$, and $25$ — followed by final testing. The results, presented in Table 2, indicate that selective weighting enhances performance across different adaptation durations and various benchmarks, confirming our hypothesis in Question 3.
Language Encoding Model.
To investigate the impact of language encoders under blurred task boundaries with paraphrased descriptions, we ablated the choice of language encoding model. Specifically, we compared our chosen Sentence Similarity (SS) Model, which excels at clustering semantically similar language descriptions, with BERT, the default language encoder from LIBERO. We selected the libero_goal benchmark for this study because its tasks are visually similar, making effective language embedding crucial for distinguishing tasks and aiding data retrieval for local adaptation.
Our experimental results yield the following observations:
(1) As illustrated in Figure 4 (a) and (b), the PCA results show that the SS Model effectively differentiates tasks, whereas BERT struggles, leading to inadequate task distinction. Consequently, as shown in Figure 4 (c), the model trained with BERT embeddings on libero_goal performs worse than the one trained with SS Model embeddings.
(2) Due to this limitation, BERT is unable to retrieve the most relevant demonstrations (those most similar to the current task from the episodic memory $\mathcal{M}$). As a result, Retrieval-based WLA with BERT does not achieve optimal performance. These two findings address Question 1.
(3) Interestingly, from Figure 4 (c), despite BERT’s low Retrieval Accuracy (RA), if it attains a moderately acceptable rate (e.g., 0.375), the local adaptation using data retrieved based on BERT embeddings can still enhance model performance during test time. This demonstrates the robustness and fault tolerance of our proposed approach, further addressing Question 5.
Table 2: Ablation Study on Selective Weighting. This table presents the performance of success rates with uniform (ULA) and weighted (WLA) local adaptation across $15$, $20$, and $25$ epochs of adaptation under three random seeds, with evaluations conducted on all $10$ tasks within the benchmarks: libero_spatial, libero_object, and libero_goal. Compared to ULA, the weighted scheme significantly improves the method’s performance on most benchmarks.
Benchmark Method 15 Epochs 20 Epochs 25 Epochs Overall ASR (%)
ASR (%) ASR (%) ASR (%)
libero_spatial ER-ULA 35.33 $\pm$ 21.21 38.17 $\pm$ 14.76 38.16 $\pm$ 17.19 37.22 $\pm$ 17.77
ER-WLA 36.16 $\pm$ 18.55 39.83 $\pm$ 19.84 37.83 $\pm$ 17.70 37.94 $\pm$ 18.57
libero_object ER-ULA 57.83 $\pm$ 25.14 60.67 $\pm$ 22.96 58.00 $\pm$ 21.84 58.83 $\pm$ 23.13
ER-WLA 58.00 $\pm$ 22.35 62.33 $\pm$ 18.70 61.50 $\pm$ 24.36 60.61 $\pm$ 21.76
libero_goal ER-ULA 61.33 $\pm$ 28.43 62.00 $\pm$ 29.61 66.17 $\pm$ 27.22 63.17 $\pm$ 28.20
ER-WLA 62.83 $\pm$ 28.15 62.33 $\pm$ 28.76 67.50 $\pm$ 28.82 64.22 $\pm$ 28.35
((a)) PCA Visualization based on BERT
((b)) PCA Visualization based on SS Model
((c)) Bar Chart of Average Success Rates and Retrieval Accuracy across $10$ tasks
Figure 4: In Figure 4(a) and Figure 4(b), PCA is used to visualize the distribution of language embeddings of 3 tasks from BERT and SS, respectively. In Figure 4(c), SS model, which distinguishes task descriptions, has higher success rate and retrieval accuracy than BERT.
6 Conclusion and Discussion
In this paper, we introduced a novel task-unaware lifelong robot learning framework that combines retrieval-based local adaptation with selective weighting during test time. Our approach enables robots to continuously learn and adapt in dynamic environments without explicit task identifiers or predefined boundaries. Leveraging an episodic memory $\mathcal{M}$, our method retrieves relevant past demonstrations based on visual and language similarities, allowing the robot to fine-tune its policy locally. The selective weighting mechanism enhances adaptation by prioritizing the most challenging segments of the retrieved demonstrations. Notably, our framework is not only robust, but is compatible with various memory-based lifelong learning methods, enhancing a robot’s ability to perform previously learned tasks as a paradigm.
A key challenge lies in the selective weighting process, particularly in finding the Separation Segment. Real-world noise, the multimodal nature of manipulation actions, and varying semantic information make it sometimes difficult to accurately identify Separation Segment in demonstration trajectories. Addressing this issue will be the focus of our future work to further improve our approach.
Appendix A Appendix
A.1 Notations
Table 3: Mathematical Notations
Symbol Description
$k$ Index of tasks, $k=1,\dots,K$
$K$ Total number of tasks
$n$ Index of retrieved demonstrations
$\tilde{N}$ Number of retrieved demonstrations
$i$ Index of samples within a demonstration
$t$ Time step
$l_{k}$ Number of samples for task $k$
$l_{n}$ Length of retrieved demonstration $n$
$\mathcal{T}_{k}$ Task $k$ (represented by multiple goal descriptions)
$\mathcal{D}_{k}$ Set of demonstrations for task $k$
$\tau^{k}_{i}$ Demonstration (trajectory) $i$ for task $k$
$\mathcal{M}$ Episodic memory buffer
$\mathbf{o}_{t}$ Observation vector at time $t$
$\mathbf{o}_{\leq t}$ Sequence of observation vectors up to time $t$ to deal with partial observability
$\mathbf{a}_{t}$ Action vector at time $t$
$\mathbf{a}_{t}^{k}$ Action vector at time $t$ for task $k$
$x_{i,n}$ Input of sample $i$ in retrieved demonstration $n$
$y_{i,n}$ Label (action) of sample $i$ in retrieved demonstration $n$
$\theta$ Model parameters
$\theta^{*}$ Optimal model parameters
$\mathbf{\theta}_{k}$ Model parameters after adaptation on task $k$
$\pi_{\theta}$ Policy parameterized by $\theta$
$\pi_{\theta}(\mathbf{s}_{\leq t},\mathcal{T}_{k})$ Policy output given states up to time $t$ and task $\mathcal{T}_{k}$
$\mathcal{L}$ Loss function
$p(y\mid x;\theta)$ Probability of label $y$ given input $x$ and parameters $\theta$
$w_{i,n}$ Weight assigned to sample $i$ in retrieved demonstration $n$ during adaptation
$\mathbb{E}$ Expectation operator
$g_{i}$ Goal descriptions in task $\mathcal{T}_{k}$
A.2 Implementation and Training Details
A.2.1 Network Architecture and Modularities
Table 4 summarizes the core components of our network architecture, while Table 5 details the input and output dimensions.
Table 4: Network architecture of the proposed Model.
Module Configuration
Pretrained Image Encoder ResNet-based R3M (Nair et al., 2022), output size: 512
Image Embedding Layer MLP, input size: 512, output size: 64
Pretrained Language Encoder Sentence Similarity (SS) Model (SentenceSimilarity, 2024), output size: 384
Language Embedding Layer MLP, input size: 384, output size: 64
Extra Modality Encoder (Proprio) MLP, input size: 9, output size: 64
Temporal Position Encoding sinusoidal positional encoding, input size: 64
Temporal Transformer heads: 6, sequence length: 10, dropout: 0.1, head output size: 64
Policy Head (GMM) modes: 5, input size: 64, output size: 7
Table 5: Inputs and Output Shape.
Modularities Shape
Image from Workspace Camera $128\times 128\times 3$
Image from Wrist Camera $128\times 128\times 3$
Max Word Length $75$
Joint States 7
Gripper States 2
Action $7$
A.2.2 Training Hyperparameters
Table 6 provides a summary of the essential hyperparameters used during training and local adaptation. The model was trained on a combination of A40, A100, and L40S GPUs, while we also leveraged multi-GPU configurations to accelerate the training process. For each task, demonstration data was initially collected and provided by LIBERO benchmark. However, due to version discrepancies that introduced visual and physical variations in the simulation, we reran the demonstrations with the latest version to obtain updated observations. It is important to note that occasional rollout failures occurred because different versions of RoboMimic Simulation (Mandlekar et al., 2021) utilize varying versions of the MuJoCo Engine (Todorov et al., 2012).
Task performance was evaluated every $10$ epochs using $20$ parallel processes to maximize efficiency. The best-performing model from these evaluations was retained for subsequent tasks. After training on each task, we reassessed the model’s performance across all previously encountered tasks.
Table 6: Hyperparameter for Training and Adaptation.
Hyperparameter Value
Batch Size $32$
Learning Rate $0.0001$
Optimizer AdamW
Betas $[0.9,0.999]$
Weight Decay $0.0001$
Gradient Clipping $100$
Loss Scaling $1.0$
Training Epochs $50$
Image Augmentation Translation, Color Jitter
Evaluation Frequency Every $10$ epochs
Number of Demos per Task Up to $50$  [^1] For each task, demonstration data was collected from LIBERO, but due to differences in simulation versions, the demonstrations were rerun in the current simulation to collect new observations, with the possibility of occasional failures during rollout (see Subsection A.2.2 for details).
Number of Demos per Task in EM $8$
Rollout Episodes before Adaptation $10$
Distance weights $[\alpha_{v},\alpha_{l}]$ for libero_spatial and libero_object $[1.0,0.5]$
Distance weights $[\alpha_{v},\alpha_{l}]$ for libero_goal $[0.5,1.0]$
Distance weights $[\alpha_{v},\alpha_{l}]$ for libero_different_scenes $[1.0,0.1]$
Weights Added for Separation Segments $0.3$
Clipping Range for Selective Weighting $2$
Default Local Adaptation Epochs $20$
A.2.3 Baseline Details
We follow the implementation of all baselines and hyperparameters for individual algorithms from (Liu et al., 2024), maintaining the same backbone model and episodic memory structure as in our approach. During the training phase, we also apply the same learning hyperparameters outlined in Table 6.
A.3 Details about Task-unaware setting
In this paper, we blur task boundaries by using multiple paraphrased descriptions that define the task goals. The following section elaborate more details about our dataset and process of task description paraphrase.
A.3.1 Datasets Structure
Our dataset inherent the dataset from LIBERO (Liu et al., 2024), maintaining all the attributes and data. Additionally, we add demo description to each demonstration to achieve task unawareness and augmented description to augment language description during training (See Figure 5). Unlike the dataset from LIBERO, which groups demonstrations together under one specific task, our dataset wrap all demonstrations with random order to eliminate the task boundary.
Figure 5: Data Structure
A.3.2 Description Paraphrase
We leverage the Phi-3-mini-4k-instruct model (mini-4k instruct, 2024) to paraphrase the task description. The process and prompts that we use are illustrated in Figure 6.
Figure 6: Paraphrase Description
A.4 Details about Selective Weighting
In this section, we introduce our Selective Weighting mechanism in detail.
A.4.1 Detailed Heuristics and Implementations
To assign weights to retrieved demonstrations, we analyze the distance between demonstration and failed rollout trajectories. Typically, the comparison distance increases as the failed rollout diverges from the demonstration.
Due to the multi-modal nature of robotic actions and visual observation noise, raw distance comparisons can be erratic. To address this, we smooth the distance curves using a moving window. Despite smoothing, the trend may remain jittery, making it challenging to pinpoint a single separation point where performance deviates. Instead, we identify a range of frames representing the Separation Segment where the distances worsen, indicating vulnerable steps in a manipulation task.
We apply two thresholds to detect the segment. Specifically, we locate frames where the distance falls between $\frac{1}{8}$ and $\frac{1}{3}$ of the maximum observed distance. We focus on the last occurrence within this range to account for possible initial divergent paths that later converge. Once identified, we extend the separation segment by $15$ frames before and after to mitigate noise effects.
For each frame within the separation segment, we add a weight of $0.3$ to the initially uniform weight vector. This process is repeated for up to five failed rollouts per retrieved demonstration. After processing all demonstrations, we clip the weights to a maximum of $2$ and normalize the weight vector to maintain a consistent loss function scale, ensuring stable gradient updates.
During adaptation, the resulting weights ($w_{t,n}$) are integrated into the loss function as described in  equation 3. This approach enhances the influence of critical samples while reducing the impact of less relevant ones, thereby improving the model’s learning efficiency.
A.4.2 Detailed Ablation Studies on Selective Weighting.
The average success rate per benchmark is illustrated in Table 2. The detailed results on each task are shown in Table 7, Table 8, and Table 9.
Table 7: Ablation Study Results on libero_object: Average Success Rates and Standard Devision for Each Task Across Epochs.
Method Task  Epoch ULA WLA
Epoch 15 Epoch 20 Epoch 25 Epoch 15 Epoch 20 Epoch 25
0 0.68 $\pm$ 0.04 0.37 $\pm$ 0.12 0.62 $\pm$ 0.04 0.57 $\pm$ 0.14 0.67 $\pm$ 0.15 0.57 $\pm$ 0.04
1 0.20 $\pm$ 0.08 0.40 $\pm$ 0.13 0.35 $\pm$ 0.22 0.35 $\pm$ 0.15 0.45 $\pm$ 0.06 0.13 $\pm$ 0.06
2 0.77 $\pm$ 0.14 0.85 $\pm$ 0.06 0.78 $\pm$ 0.15 0.90 $\pm$ 0.08 0.78 $\pm$ 0.14 0.82 $\pm$ 0.11
3 0.68 $\pm$ 0.15 0.78 $\pm$ 0.06 0.70 $\pm$ 0.03 0.70 $\pm$ 0.09 0.60 $\pm$ 0.10 0.75 $\pm$ 0.08
4 0.75 $\pm$ 0.08 0.87 $\pm$ 0.02 0.78 $\pm$ 0.07 0.70 $\pm$ 0.06 0.78 $\pm$ 0.07 0.88 $\pm$ 0.03
5 0.47 $\pm$ 0.19 0.65 $\pm$ 0.05 0.53 $\pm$ 0.04 0.37 $\pm$ 0.09 0.42 $\pm$ 0.07 0.60 $\pm$ 0.13
6 0.52 $\pm$ 0.06 0.53 $\pm$ 0.09 0.38 $\pm$ 0.16 0.65 $\pm$ 0.12 0.52 $\pm$ 0.12 0.55 $\pm$ 0.08
7 0.47 $\pm$ 0.19 0.58 $\pm$ 0.14 0.57 $\pm$ 0.09 0.58 $\pm$ 0.04 0.73 $\pm$ 0.04 0.60 $\pm$ 0.18
8 0.55 $\pm$ 0.10 0.58 $\pm$ 0.17 0.50 $\pm$ 0.13 0.58 $\pm$ 0.06 0.70 $\pm$ 0.09 0.72 $\pm$ 0.09
9 0.70 $\pm$ 0.18 0.45 $\pm$ 0.10 0.58 $\pm$ 0.03 0.40 $\pm$ 0.15 0.58 $\pm$ 0.02 0.53 $\pm$ 0.09
Table 8: Ablation Study Results on libero_goal: Average Success Rates and Standard Devision for Each Task Across Epochs.
Method Task  Epoch ULA WLA
Epoch 15 Epoch 20 Epoch 25 Epoch 15 Epoch 20 Epoch 25
0 0.62 $\pm$ 0.09 0.75 $\pm$ 0.05 0.68 $\pm$ 0.10 0.72 $\pm$ 0.04 0.83 $\pm$ 0.09 0.65 $\pm$ 0.03
1 0.88 $\pm$ 0.03 0.92 $\pm$ 0.03 0.88 $\pm$ 0.02 0.87 $\pm$ 0.06 0.88 $\pm$ 0.04 0.92 $\pm$ 0.04
2 0.65 $\pm$ 0.13 0.72 $\pm$ 0.12 0.80 $\pm$ 0.03 0.68 $\pm$ 0.08 0.70 $\pm$ 0.08 0.83 $\pm$ 0.06
3 0.38 $\pm$ 0.07 0.25 $\pm$ 0.03 0.32 $\pm$ 0.09 0.32 $\pm$ 0.12 0.38 $\pm$ 0.16 0.32 $\pm$ 0.06
4 0.88 $\pm$ 0.04 0.80 $\pm$ 0.05 0.82 $\pm$ 0.03 0.87 $\pm$ 0.04 0.77 $\pm$ 0.14 0.92 $\pm$ 0.04
5 0.60 $\pm$ 0.10 0.53 $\pm$ 0.13 0.63 $\pm$ 0.20 0.58 $\pm$ 0.07 0.62 $\pm$ 0.12 0.77 $\pm$ 0.03
6 0.15 $\pm$ 0.03 0.15 $\pm$ 0.09 0.22 $\pm$ 0.07 0.13 $\pm$ 0.04 0.22 $\pm$ 0.06 0.20 $\pm$ 0.03
7 0.93 $\pm$ 0.04 0.95 $\pm$ 0.05 1.00 $\pm$ 0.00 0.97 $\pm$ 0.02 0.93 $\pm$ 0.02 0.93 $\pm$ 0.04
8 0.78 $\pm$ 0.07 0.80 $\pm$ 0.03 0.77 $\pm$ 0.04 0.72 $\pm$ 0.06 0.67 $\pm$ 0.11 0.90 $\pm$ 0.05
9 0.25 $\pm$ 0.03 0.33 $\pm$ 0.08 0.50 $\pm$ 0.10 0.43 $\pm$ 0.17 0.23 $\pm$ 0.06 0.32 $\pm$ 0.09
Table 9: Ablation Study Results on libero_spatial: Average Success Rates and Standard Devision for Each Task Across Epochs.
Method Task  Epoch ULA WLA
Epoch 15 Epoch 20 Epoch 25 Epoch 15 Epoch 20 Epoch 25
0 0.35 $\pm$ 0.18 0.33 $\pm$ 0.07 0.47 $\pm$ 0.07 0.45 $\pm$ 0.10 0.45 $\pm$ 0.10 0.42 $\pm$ 0.13
1 0.48 $\pm$ 0.09 0.43 $\pm$ 0.04 0.48 $\pm$ 0.10 0.30 $\pm$ 0.05 0.58 $\pm$ 0.19 0.40 $\pm$ 0.13
2 0.32 $\pm$ 0.11 0.35 $\pm$ 0.13 0.28 $\pm$ 0.12 0.40 $\pm$ 0.19 0.50 $\pm$ 0.13 0.45 $\pm$ 0.13
3 0.48 $\pm$ 0.03 0.47 $\pm$ 0.09 0.60 $\pm$ 0.05 0.48 $\pm$ 0.07 0.47 $\pm$ 0.11 0.50 $\pm$ 0.00
4 0.17 $\pm$ 0.04 0.30 $\pm$ 0.03 0.13 $\pm$ 0.07 0.22 $\pm$ 0.07 0.18 $\pm$ 0.07 0.23 $\pm$ 0.02
5 0.12 $\pm$ 0.09 0.20 $\pm$ 0.08 0.28 $\pm$ 0.09 0.25 $\pm$ 0.10 0.22 $\pm$ 0.09 0.27 $\pm$ 0.02
6 0.60 $\pm$ 0.13 0.58 $\pm$ 0.02 0.47 $\pm$ 0.10 0.57 $\pm$ 0.07 0.58 $\pm$ 0.07 0.67 $\pm$ 0.03
7 0.52 $\pm$ 0.06 0.42 $\pm$ 0.02 0.38 $\pm$ 0.07 0.38 $\pm$ 0.06 0.38 $\pm$ 0.04 0.38 $\pm$ 0.06
8 0.30 $\pm$ 0.05 0.42 $\pm$ 0.08 0.30 $\pm$ 0.00 0.40 $\pm$ 0.10 0.28 $\pm$ 0.03 0.30 $\pm$ 0.03
9 0.20 $\pm$ 0.10 0.32 $\pm$ 0.09 0.42 $\pm$ 0.03 0.17 $\pm$ 0.07 0.33 $\pm$ 0.06 0.17 $\pm$ 0.04
A.5 Detailed Testing Results
We selected 20 typical scenarios among libero_90. The list of those scenarios can be found in Table 10. Additionally, the testing results of our method and baselines including ER-WLA, ER, Packnet, are listed in Table 11
Table 10: Selected Tasks for libero_different_scenes benchmark from libero_90
Task ID Initial Descriptions Scenes
1 Close the top drawer of the cabinet Kitchen scene10
2 Open the bottom drawer of the cabinet Kitchen scene1
3 Open the top drawer of the cabinet Kitchen scene2
4 Put the frying pan on the stove Kitchen scene3
5 Close the bottom drawer of the cabinet Kitchen scene4
6 Close the top drawer of the cabinet Kitchen scene5
7 Close the microwave Kitchen scene6
8 Open the microwave Kitchen scene7
9 Put the right moka pot on the stove Kitchen scene8
10 Put the frying pan on the cabinet shelf Kitchen scene9
11 Pick up the alphabet soup and put it in the basket Living Room scene1
12 Pick up the alphabet soup and put it in the basket Living Room scene2
13 Pick up the alphabet soup and put it in the tray Living Room scene3
14 Pick up the black bowl on the left and put it in the tray Living Room scene4
15 Put the red mug on the left plate Living Room scene5
16 Put the chocolate pudding to the left of the plate Living Room scene6
17 Pick up the book and place it in the front compartment of the caddy Study scene1
18 Pick up the book and place it in the back compartment of the caddy Study scene2
19 Pick up the book and place it in the front compartment of the caddy Study scene3
20 Pick up the book in the middle and place it on the cabinet shelf Study scene4
Table 11: Detailed Comparisons on libero_different_scenes Benchmark. It illustrates that after the reaching the capacity of PackNet, it could no longer deal with new tasks anymore.
Task ER-WLA ER Packnet
0 0.85 $\pm$ 0.08 0.50 $\pm$ 0.03 1.00 $\pm$ 0.00
1 0.13 $\pm$ 0.08 0.27 $\pm$ 0.06 0.83 $\pm$ 0.09
2 0.73 $\pm$ 0.09 0.72 $\pm$ 0.10 0.92 $\pm$ 0.02
3 0.40 $\pm$ 0.03 0.13 $\pm$ 0.02 0.17 $\pm$ 0.03
4 0.93 $\pm$ 0.04 0.72 $\pm$ 0.10 1.00 $\pm$ 0.00
5 1.00 $\pm$ 0.00 0.57 $\pm$ 0.16 1.00 $\pm$ 0.00
6 0.52 $\pm$ 0.04 0.52 $\pm$ 0.03 0.78 $\pm$ 0.04
7 0.82 $\pm$ 0.07 0.63 $\pm$ 0.09 0.88 $\pm$ 0.02
8 0.32 $\pm$ 0.07 0.23 $\pm$ 0.06 0.00 $\pm$ 0.00
9 0.48 $\pm$ 0.15 0.38 $\pm$ 0.12 0.00 $\pm$ 0.00
10 0.23 $\pm$ 0.06 0.03 $\pm$ 0.02 0.00 $\pm$ 0.00
11 0.20 $\pm$ 0.03 0.10 $\pm$ 0.06 0.00 $\pm$ 0.00
12 0.23 $\pm$ 0.09 0.13 $\pm$ 0.02 0.00 $\pm$ 0.00
13 0.67 $\pm$ 0.09 0.83 $\pm$ 0.04 0.00 $\pm$ 0.00
14 0.15 $\pm$ 0.03 0.13 $\pm$ 0.04 0.00 $\pm$ 0.00
15 0.68 $\pm$ 0.09 0.30 $\pm$ 0.08 0.00 $\pm$ 0.00
16 0.03 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
17 0.28 $\pm$ 0.08 0.02 $\pm$ 0.02 0.00 $\pm$ 0.00
18 0.10 $\pm$ 0.08 0.02 $\pm$ 0.02 0.00 $\pm$ 0.00
19 0.27 $\pm$ 0.16 0.58 $\pm$ 0.07 0.00 $\pm$ 0.00