### v3

# Task-free Lifelong Robot Learning with Retrieval-based Weighted Local Adaptation

Abstract
A fundamental objective in intelligent robotics is to move towards lifelong learning robot that can learn and adapt to unseen scenarios over time. However, continually learning new tasks would introduce catastrophic forgetting problems due to data distribution shifts. To mitigate this, we store a subset of data from previous tasks and utilize it in two manners: leveraging experience replay to retain learned skills and applying a novel Retrieval-based Local Adaptation technique to restore relevant knowledge. Since a lifelong learning robot must operate in task-free scenarios, where task IDs and even boundaries are not available, our method performs effectively without relying on such information. We also incorporate a selective weighting mechanism to focus on the most ”forgotten” skill segment, ensuring effective knowledge restoration. Experimental results across diverse manipulation tasks demonstrate that our framework provides a scalable paradigm for lifelong learning, enhancing robot performance in open-ended, task-free scenarios.
1 Introduction
Figure 1: Method Overview. Our approach addresses the challenge of lifelong learning without relying on task boundaries or IDs. To emulate human learning patterns, we propose a method consisting of three phases: Lifelong Learning, Reviewing, and Testing. In the Lifelong Learning phase, the robot is exposed to various demonstrations, storing a subset of the data in a storage memory $\mathcal{M}$. During the Reviewing phase before policy deployment, the method retrieves the most relevant data to locally adapt the policy network, enhancing performance in the deployment scenario.
Significant progress has been made in applying lifelong learning to domains such as computer vision (Huang et al., 2024; Du et al., 2024) and natural language processing (Shi et al., 2024; Razdaibiedina et al., 2023). However, extending lifelong learning to robotics poses additional challenges, as robots must interact with the real world under sequential decision-making constraints. Besides, the high cost and complexity of physical interactions (Zhu et al., 2022; Du et al., 2023) limit data availability, emphasizing the need for more effective strategies to mitigate catastrophic forgetting and maintain performance over robots’ lifelong run (Thrun & Mitchell, 1995). Furthermore, in realistic and scalable scenarios, lifelong learning robots must operate in a task-free setting, where they are not provided with specific task IDs or boundaries for each new task. This further complicates the challenge, as robots must continually adapt to new tasks without prior knowledge of task distinctions.
In practical lifelong robot learning settings (Liu et al., 2024, 2023), a robot learns a series of tasks in a sequential manner. Lacking future data or environments, while also having only partial or no access to past data due to storage or security restrictions, leads to significant data distribution shifts and hinders stable lifelong learning. Although numerous approaches address this catastrophic forgetting problem (Wang et al., 2024), many rely on task boundaries or explicit task identifiers to consolidate knowledge, limiting their scalability in open-ended real-world scenarios (Koh et al., 2021).
To tackle these issues, we propose a task-free memory-based approach for lifelong learning in robotic manipulation. Noteworthy, “task-free” does not mean the robot ignores the task context - it should be aware of the instructions combined with the observations to fulfill the job; rather, it means the proposed algorithm effectively handles multiple continually encountered tasks and integrates new knowledge and skills without relying on known task boundaries or explicit task IDs. The concept of “task-free” is properly defined in Section 3. Our method employs a compact storage memory $\mathcal{M}$ that holds a small set of previous tasks’ demonstrations. During training, we replay samples from $\mathcal{M}$ to preserve previously acquired knowledge and skills. Nonetheless, partial forgetting remains inevitable due to the multitasking nature of lifelong learning and limited access to past data. Inspired by human learning mechanisms — where we briefly revisit specific forgotten information to rapidly restore proficiency (Sara, 2000) — we then enable the robot to perform fast local adaptation just before policy deployment. Crucially, we use the same storage memory $\mathcal{M}$ for this adaptation, avoiding any additional storage burden.
Our system retrieves relevant demonstrations based on observation similarity (Du et al., 2023; van Dijk et al., 2024; de Masson D’Autume et al., 2019) and selectively focuses on the most challenging segments where performance degrades—typically the final steps leading to task failure. This automatic selective weighting without relying on human intervention (Spencer et al., 2022; Mandlekar et al., 2020) enables the robots to efficiently restore lost knowledge and skills, facilitating stable, task-free lifelong robot learning. Our key contributions are summarized as:
  • Task-free Retrieval-Based Local Adaptation: A novel local adaptation strategy that retrieves relevant past demonstrations from $\mathcal{M}$ to restore forgotten knowledge and skills, without requiring task IDs or boundaries.
  • Selective Weighting Mechanism: An automatic weighting scheme emphasizing the most challenging segments within the retrieved demonstrations, optimizing real-time adaptation.
  • A General Paradigm for Memory-Based Lifelong Robot Learning: Our approach serves as a plug-and-play solution, complementing existing memory-based lifelong learning algorithms and enabling skill restoration in sequences of open-ended robotics tasks.
2 Related Works
2.1 Lifelong Robot Learning
Robots operating in continually changing environments need the ability to learn and adapt on-the-fly (Thrun, 1995). In recent years, lifelong robot learning has been applied to SLAM (Yin et al., 2023; Gao et al., 2022; Vödisch et al., 2022), navigation (Kim et al., 2024), and task and motion planning scenarios (Mendez-Mendez et al., 2023). In the context of lifelong learning from demonstrations, robots can 1) acquire skills from non-technical users (Grollman & Jenkins, 2007), or 2) adapt to user-specific task preferences — Chen et al. (2023) proposed a strategy mixture approach to efficiently model new incoming demonstrations, enhancing adaptability.
Furthermore, methods have been developed to improve manipulation capabilities over a robot’s lifespan. Some approaches leverage previous data to facilitate forward transfer but suffer from catastrophic forgetting (Xie & Finn, 2022). Others maintain an expandable skill set to accommodate an increasing number of manipulation tasks (Parakh et al., 2024), or continually update models of manipulable objects for effective reuse (Lu et al., 2022). Large language models have been utilized to improve knowledge transfer (Bärmann et al., 2023; Tziafas & Kasaei, 2024; Wang et al., 2023), and hypernetworks with neural ODEs have been employed to remember long trajectories (Auddy et al., 2023) incrementally. Additionally, (Yang et al., 2022) evaluates how typical supervised lifelong learning methods can be applied in reinforcement learning scenarios for robotic tasks.
To standardize the investigation of lifelong decision-making and bridge research gaps, Liu et al. (2024) introduced LIBERO, a benchmarking platform for lifelong robot manipulation where robots learn multiple atomic manipulation tasks sequentially. Recent works exploring lifelong robot learning based on it include (Liu et al., 2023), which assigns a specific task identity to each task; (Wan et al., 2024), which requires a pre-training phase to build an initial skill set before lifelong learning; and (Lee et al., 2024), tackles multi-stage tasks by incrementally learning skill prototypes for each subgoal, which introduces additional complexities in managing subgoal sequences. However, catastrophic forgetting for lifelong robot learning remains an open challenge, especially when task IDs and boundaries are not available.
2.2 Task-free Lifelong Learning
Despite the success of lifelong learning under clearly labeled task sequences, a significant gap remains in algorithms that can operate independently of task boundaries or IDs during both training and inference, thus aligning more closely with realistic and scalable scenarios. Many approaches (Lee et al., 2020; Chen et al., 2020; Ardywibowo et al., 2022) focus on specialized parameters via expanding network architectures. Meanwhile, researchers have tackled implicit task boundaries in regularization-based methods (Aljundi et al., 2019a; Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017) by consolidating knowledge upon detecting a loss “plateau”. Additionally, (Lässig et al., 2023) introduces a bio-inspired activity-regularization approach employing selective sparsity and recurrent lateral connections, effectively enabling task-free lifelong learning without explicit task boundaries.
Memory-based algorithms further mitigate forgetting by prioritizing informative samples (Sun et al., 2022), discarding less critical examples (Koh et al., 2021), refining decision boundaries (Shim et al., 2021), or enhancing gradient diversity (Aljundi et al., 2019b). Methods aiming to exploit replay buffer in online scenarios (Mai et al., 2021; Caccia et al., 2021) have also demonstrated notable success. However, these algorithms remain largely unexplored in robotic applications that entail sequential decision-making and real-world physical interactions.
2.3 Robot Learning with Adaptation
Recent advances have shown robots adapting to dynamic environments, such as executing agile flight in strong winds (O’Connell et al., 2022), adapting quadruped locomotion through test-time search (Peng et al., 2020), and generalizing manipulation skills from limited data (Julian et al., 2020). To enable few-shot or one-shot adaptation, meta-learning has been extensively explored (Finn et al., 2017a) and successfully applied to robotics (Kaushik et al., 2020; Nagabandi et al., 2018; Finn et al., 2017b). However, meta-learning methods typically assume access to a full distribution of tasks during meta-training, with both training and testing performed on tasks sampled from this distribution. In contrast, our lifelong robot learning scenario operating sequentially lacks such access, presenting unique challenges of catastrophic forgetting.
Figure 2: Policy Backbone Architecture used in Training and Testing. We input various data modalities into the system, including demonstration images, language descriptions, and the robot arm’s proprioceptive input (joint and gripper states). Pretrained R3M (Nair et al., 2022) and SentenceSimilarity (2024) models process the image and language data respectively. Along with the proprioceptive states processed by an MLP, the embeddings are concatenated and passed through a Transformer to generate temporal embeddings. A GMM (Gaussian Mixture Model) is then used as the policy head to sample actions for the robot. Throughout both training and testing, we utilize a storage memory to store a subset of demonstrations gathered throughout the training process.
3 Preliminaries
To model realistic settings for lifelong robot learning, we define a set of manipulation tasks as $\mathbb{T}=\{\mathcal{T}_{k}\},k=1,2,\dots,T$, where each $\mathcal{T}_{k}$ encompasses a distribution over environmental variations $E_{k}$ (e.g., object positions, robot initial states) and language descriptions $G_{k}$ to guide robot’s actions (e.g., “pick the bottle and put it into the basket,” “place the bottle in the basket please”). From each $\mathcal{T}_{k}$, we sample specific environmental settings $e\sim E_{k}$ and language descriptions $g\sim G_{k}$ to generate a concrete scenario $\mathcal{S}_{n}^{k}\sim p(\mathcal{T}_{k})$, which also serves as the basis for collecting demonstrations $\tau_{n}^{k}$. Multiple demonstrations form the training dataset $\mathcal{D}_{k}=\{\tau^{k}_{n}\},n=1,2,\dots,N$ for task $\mathcal{T}_{k}$.
Notably, multiple tasks may share overlapping distributions in either environmental settings or language descriptions. This natural setting closely mirrors real-world conditions, where it is difficult to determine which task generated a given scenario - tasks are not always divisible. This ambiguity underpins the proposed method’s task-free design, which emphasizes retrieving relevant information rather than relying on task boundaries or IDs.
The robot utilizes a visuomotor policy learned through behavior cloning to execute manipulation tasks by mapping sensory inputs and language description to motor actions. The policy is trained by minimizing the discrepancy between the predicted actions and expert actions from demonstrations. Specifically, we optimize the following loss function across a sequence of tasks $\mathbb{T}$ with $\mathcal{D}_{k}$. Notably, $\mathcal{D}_{k}$ is only partially accessible for $k<K$ from the storage memory $\mathcal{M}$, where $K$ denotes the current task:
$\theta^{*}$ $=\arg\min_{\theta}\;\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(o_{t},a_{t})\sim\mathcal{D}_{k},\;g\sim G_{k}}$ (1)
$\quad\Biggl{[}\sum_{t=0}^{l_{k}}\mathcal{L}\bigl{(}\pi_{\theta}(o_{\leq t},g),\,a_{t}\bigr{)}\Biggr{]}$
$\theta$ denotes the model parameters, $l_{k}$ represents the number of samples for task $k$, $o_{\leq t}$ denotes the sequence of observations up to time $t$ in demonstration $n$ (i.e., $o_{\leq t}=(o_{0},o_{1},\ldots,o_{t})$), and $a_{t}$ is the expert action at time $t$. The policy output, $\pi_{\theta}(o_{\leq t},g)$, is conditioned on both the observation sequence and the language description.
By optimizing this objective function, the policy effectively continues learning new knowledge and skills in its life span, without the need for task boundaries or IDs, thereby facilitating robust and adaptable task-free lifelong learning.
Figure 3: Data Retrieval. Episodic Memory $\mathcal{M}$ randomly stores a few demonstrations collected during lifelong learning. To retrieve a small number of demonstrations most similar to the current scenario, we compute a weighted distance (Eq (4.1)) using both image and language embeddings. In the storage memory $\mathcal{M}$, red and green circles denote relevant and irrelevant demonstrations, respectively, which include language descriptions, visual observations, expert actions, and joint and gripper states. The retrieved demonstrations are then used for Weighted Local Adaptation.
4 Retrieval-based Weighted Local Adaptation (RWLA) for Lifelong Robot Learning
In this section, we outline our proposed method - Retrieval-based Weighted Local Adaptation (RWLA) - depicted in Figure 1, with corresponding pseudocode in Algorithm 1. To effectively interact with complex physical environments, the network integrates multiple input modalities, including visual inputs from workspace and wrist cameras, proprioceptive inputs of joint and gripper states, and language descriptions.
Instead of training all modules jointly in an end-to-end manner, we employ pretrained visual and language encoders that leverage prior semantic knowledge. Pretrained encoders enhance performance on downstream manipulation tasks (Liu et al., 2023) and are well-suited to differentiate between various scenarios and tasks. They produce consistent representations that are essential both for managing multiple tasks along the lifelong training and for retrieving relevant data to support our proposed local adaptation before policy deployment.
When learning new tasks, the robot preserves previously acquired skills by replaying prior manipulation demonstrations stored in storage memory $\mathcal{M}$ (Chaudhry et al., 2019). Trained with the combined data from the latest scenarios and $\mathcal{M}$, the model can acquire new skills while mitigating catastrophic forgetting of old tasks, thereby maintaining a balance between stability and plasticity (Wang et al., 2024). Figure 2 illustrates the network architecture, and implementation details are provided in Appendix A.2.
4.1 Data Retrieval
The proposed task-free lifelong learning algorithm retrieves relevant demonstrations from $\mathcal{M}$ based on similarity to the deployment scenario $S_{deploy}$. Besides, due to the blurry task boundaries, some tasks share similar visual observations but differ in their task objectives, while others have similar goals but involve different backgrounds, objects, etc. To account for these variations, the retrieval process compares both visual inputs from the workspace camera (Du et al., 2023) and language descriptions (de Masson D’Autume et al., 2019) using $L_{2}$ distances of their embeddings, following a simple rule:
$\mathcal{D}_{R}=\alpha_{v}\cdot\mathcal{D}_{v}+\alpha_{l}\cdot\mathcal{D}_{l},$
where $\mathcal{D}_{R}$ is the weighted retrieval distance, $\mathcal{D}_{v}$ represents the distance between the embeddings of the scene observation from the workspace camera, and $\mathcal{D}_{l}$ depicts the distance between the language description embeddings. The parameters $\alpha_{v}$ and $\alpha_{l}$ control the relative importance of visual and language-based distances. Based on the distances $\mathcal{D}_{R}$, the most relevant demonstrations can be retrieved from $\mathcal{M}$, as illustrated from Figure 3.
4.2 Weighted Local Adaptation
Figure 4: Trajectory and Weighting Visualizations. To identify the point of failure, we compute the similarity between the retrieved demonstrations and failed trajectories at each frame. Once the separation segment is detected, higher weights are assigned to the frames in the segment of retrieved demonstrations during local adaptation.
4.2.1 Learn from Errors by Selective Weighting
To make the best use of the limited data, we enhance their utility by assigning weights to critical or vulnerable segments in each retrieved demonstration. Specifically, before local adaptation, the robot performs several rollouts on the encountered scenario using the current model trained during the Lifelong Learning phase. This procedure allows us to evaluate the model’s performance and identify any forgetting effects (as illustrated in step 2, the Reviewing phase in Figure 1).
If a trial fails, we compare each image in the retrieved demonstrations against all images from the failed trajectories using $L_{2}$ distance of their embeddings. This comparison yields an Embedding Distance Matrix (EDM, shown in Figure 10) for each retrieved demonstration, where each value represents an embedding distance of a demonstration frame and an image from the failed rollout. This metric determines whether a particular frame has occurred during a failed rollout. Through this process, we identify the Separation Segment — frames in a demonstration where the failed rollout’s behavior starts to diverge from the demonstration (see Figure 4). Since these Separation Segments highlight expected behaviors that did not occur, we consider them critical points contributing to failure. We assign higher weights to these frames which will scale the losses during local adaptation. Detailed heuristics and implementation specifics are provided in Appendix A.4.
4.2.2 Local Adaptation with Fast Finetuning
Finally, we fine-tune the network’s parameters to better adapt to the deploying scenario $S_{deploy}$ using the retrieved demonstrations from $\mathcal{M}$, focusing more on the Separation Segments identified through selective weighting. Despite this limited data, our experiments demonstrate that the model can effectively recover learned knowledge and skills and improve robot’s performance across various tasks. Overall, the proposed weighted local adaptation is formalized as follows:
$\mathbf{\theta}^{*}=\arg\min_{\mathbf{\theta}}\sum_{n=1}^{\tilde{N}}\sum_{t=1}^{l_{n}}w_{t,n}\mathcal{L}\left(\pi_{\theta}(o_{\leq t,n},g_{n}),a_{t,n}\right)$ (2)
where $\tilde{N}$ is the number of retrieved demonstrations, $l_{n}$ is the length of demonstration $n$, and $w_{t,n}$ is the weight assigned to sample $t$ in demonstration $n$. The variables $o_{\leq t,n}$ and $a_{t,n}$ denote the sequence of observations up to time $t$ and the corresponding expert action, respectively, while $g_{n}$ is the language description for demonstration $n$. The parameter $\mathbf{\theta}$ represents the network’s parameters before local adaptation.
5 Experiments
We conduct a comprehensive set of experiments to evaluate the effectiveness of RWLA for task-free lifelong robot learning. Specifically, our experiments aim to address the following key questions:
  1. 1.
    Effect of Blurry Task Boundaries: How do the blurry task boundaries influence the model’s performance and data retrieval during testing?
  2. 2.
    Advantages of RWLA: Does the proposed approach enhance the robot’s performance across diverse tasks?
  3. 3.
    Impact of Selective Weighting: Is selective weighting based on rollout failures effective?
  4. 4.
    Generalizability: Can our method be applied to different memory-based lifelong robot learning approaches, serving as a paradigm that enhances performance?
  5. 5.
    Robustness: How robust is our approach to imperfect demonstration retrieval, particularly when ambiguous task boundaries cause retrieved examples to mismatch the deploying scenario?
5.1 Experimental Setup
5.1.1 Benchmarks
We evaluate our proposed methods using LIBERO benchmarks (Liu et al., 2024): libero_spatial, libero_object, libero_goal, and libero_different_scenes. These environments feature a variety of task goals, objects, and layouts. The first three benchmarks all include $10$ distinct task goals (e.g., “Put the bottle into the basket.”, “Open the middle drawer of the cabinet.”), each with up to $50$ demonstrations collected from sampled simulation scenarios with different initial states of objects and the robot. Specifically, libero_different_scenes is created from LIBERO’s LIBERO_90, which encompasses $20$ tasks from distinct scenes.
We paraphrased the assigned single task goal into diverse language descriptions to obscure task boundaries (See Figure 7). These enriched language descriptions were generated by rephrasing the original task goal from the benchmark using a large language model provided by Phi-3-mini-4k-instruct Model (mini-4k instruct, 2024), ensuring consistent meanings while varying phraseology and syntax. Please see Section A.3 for more details.
5.1.2 Baselines
We evaluate our proposed method against the following baseline approaches:
  1. 1.
    Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017): A regularization-based approach that relies on task boundaries and restricts network parameters’ updates to prevent catastrophic forgetting of previously learned tasks.
  2. 2.
    Experience Replay (ER) (Chaudhry et al., 2019): A core component of our training setup, ER utilizes a storage memory to replay past demonstrations, helping the model maintain previously acquired skills and mitigate forgetting.
  3. 3.
    Average Gradient Episodic Memory (AGEM) (Hu et al., 2020): Employs a memory buffer to constrain gradients during the training of new tasks, ensuring that updates do not interfere with performance on earlier tasks.
  4. 4.
    AGEM-RWLA: An extension of AGEM that incorporates RWLA before policy deployment, enhancing the model’s ability to adapt to specific scenarios. This allows us to assess the generalizability of our proposed method as a paradigm framework on other memory-based lifelong learning approaches.
  5. 5.
    PackNet (Mallya & Lazebnik, 2018): An architecture-based lifelong learning algorithm that iteratively prunes the network after training each task, preserving essential nodes while removing less critical connections to accommodate subsequent tasks. However, its pruning and post-training phases rely heavily on clearly defined task IDs, making PackNet a reference baseline when the IDs are well-defined.
Table 1: Comparison with Baselines. The Average Success Rates (ASR, %) and standard deviations (STD, %) across various baselines are shown below. We provide PackNet’s performance as a reference point for cases where task IDs are accessible. Both EWC and vanilla AGEM demonstrate weak performance across all benchmarks. Under our Retrieval-based Weighted Local Adaptation (RWLA) paradigm, both ER and AGEM show significant improvements over their vanilla counterparts, highlighting the effectiveness of RWLA.
[…]
EWC PackNet AGEM AGEM-RWLA ER ER-RWLA (ours)
libero_spatial 0.0 $\pm$ 0.0 53.17 $\pm$ 7.25 7.33 $\pm$ 1.61 35.83 $\pm$ 9.39 15.67 $\pm$ 6.66 39.83 $\pm$ 8.13
libero_object 1.50 $\pm$ 0.50 73.67 $\pm$ 1.89 27.17 $\pm$ 4.65 51.17 $\pm$ 3.62 56.50 $\pm$ 3.97 62.33 $\pm$ 1.76
libero_goal 0.33 $\pm$ 0.58 66.33 $\pm$ 4.19 10.83 $\pm$ 3.55 58.67 $\pm$ 3.21 52.33 $\pm$ 5.20 62.33 $\pm$ 4.93
libero_different_scenes 2.58 $\pm$ 1.14 32.92 $\pm$ 0.63 20.43 $\pm$ 1.17 41.75 $\pm$ 5.66 34.08 $\pm$ 2.47 45.17 $\pm$ 0.38
Table 2: Ablation Study on Selective Weighting. This table presents ASR (%) and their STD (%) for uniform (RULA) and weighted (RWLA) local adaptation across $15$, $20$, and $25$ epochs of adaptation under three random seeds, with evaluations conducted on all $10$ tasks within the benchmarks: libero_spatial, libero_object, and libero_goal. Compared to RULA, selective weighting scheme improves the method’s performance on most benchmarks.
Benchmark Method 15 Epochs 20 Epochs 25 Epochs Overall ASR
libero_spatial RULA 35.33 $\pm$ 9.07 38.17 $\pm$ 5.06 38.17 $\pm$ 3.33 37.22 $\pm$ 5.64
RWLA 36.17 $\pm$ 8.33 39.83 $\pm$ 8.13 37.83 $\pm$ 3.75 37.94 $\pm$ 6.32
libero_object RULA 57.83 $\pm$ 9.07 60.67 $\pm$ 3.01 58.00 $\pm$ 4.92 58.83 $\pm$ 5.55
RWLA 58.00 $\pm$ 4.44 62.33 $\pm$ 1.76 61.50 $\pm$ 1.32 60.61 $\pm$ 3.18
libero_goal RULA 61.33 $\pm$ 3.69 62.00 $\pm$ 2.60 66.17 $\pm$ 5.58 63.17 $\pm$ 4.24
RWLA 62.83 $\pm$ 7.65 62.33 $\pm$ 4.93 67.50 $\pm$ 1.80 64.22 $\pm$ 5.26
5.1.3 Metrics
We focus on the success rate of task execution, as it is a crucial metric for manipulation tasks in interactive robotics. Consequently, we adopt the Average Success Rate (ASR) as our primary evaluation metric to address the challenge of catastrophic forgetting within the lifelong learning framework, evaluating success rates on three random seeds across all diverse tasks within each benchmark.
5.1.4 Model, Training, and Evaluation
As illustrated in Figure 2, our model utilizes pretrained encoders for visual and language inputs: R3M (Nair et al., 2022) for visual encoding, Sentence Similarity model (SS Model) (SentenceSimilarity, 2024) for language embeddings, and a trainable MLP-based network to encode proprioceptive inputs. Embeddings from ten consecutive time steps are processed through a transformer-based temporal encoder, with the resulting output passed to a GMM-based policy head for action sampling. Specifically, R3M, a ResNet-based model trained on egocentric videos using contrastive learning, captures temporal dynamics and semantic features from scenes, while Sentence Similarity Model captures semantic meanings in language descriptions, enabling the model to differentiate between various instructions.
To standardize the comparisons with baseline lifelong robot learning algorithms in LIBERO benchmarks, the model first undergoes a Lifelong learning phase, where it is trained sequentially on demonstrations from $10$ or $20$ tasks, depending on the specific benchmark, with each task trained for $50$ epochs. A small number of demonstrations from each task is stored in $\mathcal{M}$, playing a dual-use for experience replay and RWLA. Every $10$ epochs, we check the model’s performance and save the version that achieves the highest Success Rate to prevent over-fitting.
After training on all tasks sequentially, we conduct reviewing and testing on various scenarios sampled from each task for comprehensive analysis. During the reviewing stage, we firstly evaluate potential forgetting by having the agent perform $10$ rollout episodes on the deployment scenario $\mathcal{S}_{deploy}$. We then retrieve the most similar demonstrations from $\mathcal{M}$ and fine-tune the model for only $20$ epochs using the retrieved demonstrations with selective weighting. Finally, we deploy the adapted model for $20$ episodes—the testing phase—to assess performance improvements. All training, local adaptation, and testing in the benchmarks are conducted using three random seeds ($1$, $21$, and $42$) to reduce the impact of randomness.
5.2 Results
5.2.1 Comparison with Baselines
To address Question 2, we compared RWLA, with all baseline approaches. As shown in Table 1, ER-RWLA consistently outperforms baselines of EWC, AGEM, ER, and AGEM-RWLA. By incorporating local adaptation before policy deployment — our method mirrors how humans review and reinforce knowledge when it is partially forgotten — the continually learning robot could also regain its proficiency on previous tasks.
In contrast, PackNet serves as a reference method, as it requires explicit task IDs. Noteworthy, as the number of tasks increases, the network’s trainable capacity under PackNet diminishes, leaving less flexibility for future tasks. This limitation becomes evident in the libero_different_scenes benchmark, which includes 20 tasks, see Table 8. PackNet’s success rate drops significantly for later tasks, resulting in poor overall performance and highlighting its constraints on plasticity compared with the proposed ER-RWLA.
Additionally, when we applied RWLA to the AGEM baseline (AGEM-RWLA), it also improved its performance, demonstrating the effectiveness of our method as a paradigm for memory-based lifelong robot learning methods. These findings support our conclusions regarding Question 4.
(a) PCA based on BERT
(b) PCA based on SS Model
(c) Bar Chart of Average Success Rates and Retrieval Accuracy across $10$ tasks
Figure 5: In Figures 5(a) and 5(b), PCA is used to visualize the distribution of language embeddings of 3 tasks from BERT and SS, respectively. In Figure 5(c), the SS model, which distinguishes task descriptions, has higher average success rates (ASR) and retrieval accuracy (RA) than BERT. The error bars represent the standard deviations of ASR and RA for each task over 20 repetitions with 3 random seeds.
5.2.2 Ablation Studies
We performed two ablation studies to validate the effectiveness of our implementation choices and address Questions 1, 3, and 5.
Selective Weighting.
In the first ablation, we evaluated the impact of selective weighting on libero_spatial, libero_object, and libero_goal benchmarks to demonstrate its importance for effective local adaptation. We compared a variant of RWLA: RULA, which applies uniform local adaptation without selective weighting, adapting retrieved demonstrations uniformly. Both methods are trained with ER.
Since early stopping during local adaptation at test time is infeasible, and training can be unstable, particularly regarding manipulation success rates, we conducted RWLA using three different numbers of epochs — $15$, $20$, and $25$. The results presented in Table 2, indicate that selective weighting enhances performance across different adaptation durations and various benchmarks, addressing Question 3.
Language Encoding Model.
To investigate the impact of language encoders under blurred task boundaries with paraphrased descriptions, we ablated the choice of language encoding model. Specifically, we compared our chosen Sentence Similarity (SS) Model, which excels at clustering semantically similar language descriptions, with BERT, the default language encoder from LIBERO. We selected the libero_goal benchmark for this study because its tasks are visually similar, making effective language embedding crucial for distinguishing tasks and aiding data retrieval for local adaptation.
Our experimental results yield the following observations:
(1) As illustrated in Figure 5 (a) and (b), the PCA results show that the SS Model effectively differentiates tasks, whereas BERT struggles, leading to inadequate task distinction. Consequently, as shown in Figure 5 (c), the model trained with BERT embeddings on libero_goal performs worse than the one trained with SS Model embeddings.
(2) Due to this limitation, BERT is unable to retrieve the most relevant demonstrations (those most similar to the current scenario from the storage memory $\mathcal{M}$). As a result, RWLA with BERT does not achieve optimal performance. These two findings address Question 1.
(3) Interestingly, from Figure 5 (c), despite BERT’s low Retrieval Accuracy (RA), if it attains a moderately acceptable rate (e.g., 0.375), the RWLA based on BERT embeddings can still enhance model performance. This demonstrates the robustness and fault tolerance of our proposed approach, further addressing Question 4 and 5.
6 Conclusion and Discussion
In this paper, we introduced a novel task-free lifelong robot learning framework that combines retrieval-based local adaptation with selective weighting before policy deployment. Our approach enables robots to continuously learn and restore forgotten knowledge and skills in dynamic environments without relying on task boundaries or IDs. Notably, our framework is not only robust, but is compatible with various memory-based lifelong learning methods, enhancing a robot’s ability to perform previously learned tasks as a paradigm.
A limitation of our framework is the scalability of the storage memory $\mathcal{M}$, as we continuously accumulate demonstrations. However, since image embeddings—serving dual purposes (input to the manipulation policy and data retrieval for local adaptation)—are generated by a pre-trained model, our approach is naturally extendable: this allows for significant storage reduction in future implementations, by simply storing smaller embeddings instead of raw images in $\mathcal{M}$.
Appendix A Appendix
A.1 Notations
Table 3: Mathematical Notations
Symbol Description
$k$ Index of tasks, $k=1,\dots,K$
$K$ Total number of tasks
$n$ Index of retrieved demonstrations
$\tilde{N}$ Number of retrieved demonstrations
$i$ Index of samples within a demonstration
$t$ Time step
$l_{k}$ Number of samples for task $k$
$l_{n}$ Length of retrieved demonstration $n$
$\mathcal{T}_{k}$ Task $k$ (represented by multiple goal descriptions)
$\mathcal{D}_{k}$ Set of demonstrations for task $k$
$\tau^{k}_{i}$ Demonstration (trajectory) $i$ for task $k$
$\mathcal{M}$ Episodic memory buffer
$\mathbf{o}_{t}$ Observation vector at time $t$
$\mathbf{o}_{\leq t}$ Sequence of observation vectors up to time $t$ to deal with partial observability
$\mathbf{a}_{t}$ Action vector at time $t$
$\mathbf{a}_{t}^{k}$ Action vector at time $t$ for task $k$
$x_{i,n}$ Input of sample $i$ in retrieved demonstration $n$
$y_{i,n}$ Label (action) of sample $i$ in retrieved demonstration $n$
$\theta$ Model parameters
$\theta^{*}$ Optimal model parameters
$\mathbf{\theta}_{k}$ Model parameters after adaptation on task $k$
$\pi_{\theta}$ Policy parameterized by $\theta$
$\pi_{\theta}(\mathbf{s}_{\leq t},\mathcal{T}_{k})$ Policy output given states up to time $t$ and task $\mathcal{T}_{k}$
$\mathcal{L}$ Loss function
$p(y\mid x;\theta)$ Probability of label $y$ given input $x$ and parameters $\theta$
$w_{i,n}$ Weight assigned to sample $i$ in retrieved demonstration $n$ during adaptation
$\mathbb{E}$ Expectation operator
$g_{i}$ Goal descriptions in task $\mathcal{T}_{k}$
Algorithm 1 RWLA for Task-free Lifelong Robot Learning
Lifelong Learning Phase:
  1. 1.
    Initialize model parameter $\theta$, storage memory ${}\mathcal{M}=\{\}$, and tasks $\{\mathcal{T}_{i}\}$, $i=1,2,\dots,T$
  2. 2.
    $K\in\{1,2,\dots,T\}$
    1. (a)
      Train $\theta$ on $\mathcal{D}_{K}\cup\mathcal{M}$ using Eq 1
    2. (b)
      Randomly store a small number of demonstrations from $\mathcal{D}_{K}$ into $\mathcal{M}$
During deployment, robot encounters a testing scenario $\mathcal{S}_{deploy}\sim p(\mathcal{T}_{i}),1\leq i\leq T$:
Reviewing Phase:
  1. 1.
    Rollout $10$ episodes on $\mathcal{S}_{deploy}$ to assess robot’s performance with $\theta$
  2. 2.
    Retrieve $\tilde{N}$ demonstrations from $\mathcal{M}$ based on embedding distance using Eq 4.1 (4.1)
  3. 3.
    Compute $w_{t,n}$ based on selective weighting (4.2.1)
  4. 4.
    $\theta^{\prime}\leftarrow$ Locally adapt $\theta$ using Eq 2 as skill restoration within limited epochs (4.2.2)
Testing Phase: Test $\theta^{\prime}$ in $\mathcal{S}_{deploy}$
A.2 Implementation and Training Details
A.2.1 Network Architecture and Modularities
Table 4 summarizes the core components of our network architecture, while Table 5 details the input and output dimensions.
Table 4: Network architecture of the proposed Model.
Module Configuration
Pretrained Image Encoder ResNet-based R3M (Nair et al., 2022), output size: 512
Image Embedding Layer MLP, input size: 512, output size: 64
Pretrained Language Encoder Sentence Similarity (SS) Model (SentenceSimilarity, 2024), output size: 384
Language Embedding Layer MLP, input size: 384, output size: 64
Extra Modality Encoder (Proprio) MLP, input size: 9, output size: 64
Temporal Position Encoding sinusoidal positional encoding, input size: 64
Temporal Transformer heads: 6, sequence length: 10, dropout: 0.1, head output size: 64
Policy Head (GMM) modes: 5, input size: 64, output size: 7
Table 5: Inputs and Output Shape.
Modularities Shape
Image from Workspace Camera $128\times 128\times 3$
Image from Wrist Camera $128\times 128\times 3$
Max Word Length $75$
Joint States 7
Gripper States 2
Action $7$
A.2.2 Training Hyperparameters
Table 6 provides a summary of the essential hyperparameters used during training and local adaptation. The model training was conducted using a combination of A40, A100, and L40S GPUs in a multi-GPU configuration to optimize the training process. This distributed computing setup significantly enhanced efficiency, reducing the training time per benchmark from 12 hours on a single GPU to 6 hours using 3 GPUs in parallel. For each task, demonstration data was initially collected and provided by LIBERO benchmark. However, due to version discrepancies that introduced visual and physical variations in the simulation, we reran the demonstrations with the latest version to obtain updated observations. It is important to note that occasional rollout failures occurred because different versions of RoboMimic Simulation (Mandlekar et al., 2021) utilize varying versions of the MuJoCo Engine (Todorov et al., 2012).
Task performance was evaluated every $10$ epochs using $20$ parallel processes to maximize efficiency. The best-performing model from these evaluations was retained for subsequent tasks. After training on each task, we reassessed the model’s performance across all previously encountered tasks.
Table 6: Hyperparameter for Training and Adaptation.
Hyperparameter Value
Batch Size $32$
Learning Rate $0.0001$
Optimizer AdamW
Betas $[0.9,0.999]$
Weight Decay $0.0001$
Gradient Clipping $100$
Loss Scaling $1.0$
Training Epochs $50$
Image Augmentation Translation, Color Jitter
Evaluation Frequency Every $10$ epochs
Number of Demos per Task Up to $50$  [^1] For each task, demonstration data was collected from LIBERO, but due to differences in simulation versions, the demonstrations were rerun in the current simulation to collect new observations, with the possibility of occasional failures during rollout (see Subsection A.2.2 for details).
Number of Demos per Task in $\mathcal{M}$ ($\tilde{N}$) $8$
Rollout Episodes before Adaptation $10$
Distance weights $[\alpha_{v},\alpha_{l}]$ for libero_spatial and libero_object $[1.0,0.5]$
Distance weights $[\alpha_{v},\alpha_{l}]$ for libero_goal $[0.5,1.0]$
Distance weights $[\alpha_{v},\alpha_{l}]$ for libero_different_scenes $[1.0,0.1]$
Weights Added for Separation Segments $0.3$
Clipping Range for Selective Weighting $2$
Default Local Adaptation Epochs $20$
A.2.3 Baseline Details
We follow the implementation of baselines and hyperparameters for individual algorithms from (Liu et al., 2024), maintaining the same backbone model and storage memory structure as in our approach. During training, we also apply the same learning hyperparameters outlined in Table 6.
A.3 Details about Blurred Task Boundary setting
In this paper, we blur task boundaries by using multiple paraphrased descriptions that define the task goals. The following section elaborate more details about our dataset and process of task description paraphrase.
A.3.1 Datasets Structure
Our dataset inherent the dataset from LIBERO (Liu et al., 2024), maintaining all the attributes and data. Additionally, we add demo description to each demonstration to blur task boundary and augment language description during training (See Figure 6). Unlike the dataset from LIBERO, which groups demonstrations together under one specific task, our dataset wrap all demonstrations with random order to eliminate the task boundary.
Figure 6: Data Structure
A.3.2 Description Paraphrase
We leverage the Phi-3-mini-4k-instruct model (mini-4k instruct, 2024) to paraphrase the task description. The process and prompts that we use are illustrated in Figure 7. As shown for the libero_spatial task in Figure 8, both BERT and Sentence Similarity Model struggle to distinguish tasks based on embeddings from the paraphrased descriptions. This observation further underscores the task-blurry setting in our experiments.
Figure 7: Paraphrase Description
(a) PCA Visualization based on BERT
(b) PCA Visualization based on SS Model
Figure 8: Task Blurry Effect on libero_spatial benchmark. After paraphrasing the task descriptions, both BERT and SS models struggle to distinguish the tasks in libero_spatial.
Figure 9: Hyperparameter Sensitivity Check.
A.4 Details about Selective Weighting
In this section, we introduce our Selective Weighting mechanism in detail.
A.4.1 Detailed Heuristics and Implementations
To assign weights to retrieved demonstrations, we analyze the image embedding distance between demonstration and failed rollout trajectories. Typically, the embedding distance increases as the failed rollout diverges from the demonstration. We selectively add weights for the frames in the retrieved demonstration using the Embedding Distance Curve (EDC), derived from the Embedding Distance Matrix (EDM), as illustrated in Figure 10.
Due to the multi-modal nature of robotic actions and visual observation noise, raw embedding distances can be erratic. To mitigate this, we smooth the EDC using a moving average window. Despite smoothing, the curve may remain jittery, making it difficult to pinpoint a single divergence point. Therefore, we identify a range of frames, termed the Separation Segment, where the distances increase, indicating vulnerable steps that lead to task failure.
We apply two thresholds to identify the segment: a lower threshold at $\frac{1}{8}$ and an upper threshold at $\frac{1}{3}$ of the maximum observed distance in EDC. We locate frames where the smoothed embedding distance falls within this range, focusing on the last occurrence to account for initial divergences that may later converge. We then extend this segment by $15$ frames before and after to mitigate noise effects.
For each frame within the Separation Segment, we increment the corresponding weight in the initially uniform weight vector by $0.3$. This process is repeated for up to five failed rollouts per retrieved demonstration. After processing all demonstrations, we clip the weights to a maximum of $2$ and normalize the weight vector to maintain consistent loss scaling and ensure stable gradient updates. During adaptation, the resulting weights ($w_{t,n}$) are integrated into the loss function as described in Equation (2). This selective weighting emphasizes critical samples while reducing the influence of less relevant ones, thereby enhancing the model’s learning efficiency.
A.4.2 Hyperparameter Sensitivity Analysis on Selective Weighting.
Figure 9 presents the sensitivity analysis of the hyperparameters—lower threshold ($\frac{1}{8}$), higher threshold ($\frac{1}{3}$), and padding step ($15$ steps)—used to identify Separation Segments during selective weighting. The experiments are conducted on three random seeds as well. The results demonstrate that our proposed method’s performance is robust to variations in these hyperparameters.
A.5 Detailed Testing Results
We selected 20 typical scenarios among libero_90. The list of those scenarios can be found in Table 7. Additionally, the testing results of our method and baselines including ER-RWLA, ER, Packnet, are listed in Table 8
A.6 Discussion on Potential Forgetting during Local Adaptation
Our method addresses this issue through a robust deployment strategy. After sequential learning, we preserve the final model as a stable foundation. For each testing scenario, we fine-tune a copy of this model using our weighted local adaptation mechanism. Crucially, we always return to the preserved final model for subsequent scenarios, ensuring that each adaptation starts from the same well-trained baseline and previous adaptations do not influence future ones. This approach keeps local adaptations isolated and prevents the accumulation of forgetting effects.
Table 7: Selected Tasks for libero_different_scenes benchmark from libero_90
Task ID Initial Descriptions Scenes
1 Close the top drawer of the cabinet Kitchen scene10
2 Open the bottom drawer of the cabinet Kitchen scene1
3 Open the top drawer of the cabinet Kitchen scene2
4 Put the frying pan on the stove Kitchen scene3
5 Close the bottom drawer of the cabinet Kitchen scene4
6 Close the top drawer of the cabinet Kitchen scene5
7 Close the microwave Kitchen scene6
8 Open the microwave Kitchen scene7
9 Put the right moka pot on the stove Kitchen scene8
10 Put the frying pan on the cabinet shelf Kitchen scene9
11 Pick up the alphabet soup and put it in the basket Living Room scene1
12 Pick up the alphabet soup and put it in the basket Living Room scene2
13 Pick up the alphabet soup and put it in the tray Living Room scene3
14 Pick up the black bowl on the left and put it in the tray Living Room scene4
15 Put the red mug on the left plate Living Room scene5
16 Put the chocolate pudding to the left of the plate Living Room scene6
17 Pick up the book and place it in the front compartment of the caddy Study scene1
18 Pick up the book and place it in the back compartment of the caddy Study scene2
19 Pick up the book and place it in the front compartment of the caddy Study scene3
20 Pick up the book in the middle and place it on the cabinet shelf Study scene4
Table 8: Detailed Comparisons on libero_different_scenes Benchmark. It illustrates that after reaching the capacity of PackNet, its performance on new tasks would drop drastically.
Task ER-RWLA ER Packnet
0 0.85 $\pm$ 0.08 0.50 $\pm$ 0.03 1.00 $\pm$ 0.00
1 0.13 $\pm$ 0.08 0.27 $\pm$ 0.06 0.83 $\pm$ 0.09
2 0.73 $\pm$ 0.09 0.72 $\pm$ 0.10 0.92 $\pm$ 0.02
3 0.40 $\pm$ 0.03 0.13 $\pm$ 0.02 0.17 $\pm$ 0.03
4 0.93 $\pm$ 0.04 0.72 $\pm$ 0.10 1.00 $\pm$ 0.00
5 1.00 $\pm$ 0.00 0.57 $\pm$ 0.16 1.00 $\pm$ 0.00
6 0.52 $\pm$ 0.04 0.52 $\pm$ 0.03 0.78 $\pm$ 0.04
7 0.82 $\pm$ 0.07 0.63 $\pm$ 0.09 0.88 $\pm$ 0.02
8 0.32 $\pm$ 0.07 0.23 $\pm$ 0.06 0.00 $\pm$ 0.00
9 0.48 $\pm$ 0.15 0.38 $\pm$ 0.12 0.00 $\pm$ 0.00
10 0.23 $\pm$ 0.06 0.03 $\pm$ 0.02 0.00 $\pm$ 0.00
11 0.20 $\pm$ 0.03 0.10 $\pm$ 0.06 0.00 $\pm$ 0.00
12 0.23 $\pm$ 0.09 0.13 $\pm$ 0.02 0.00 $\pm$ 0.00
13 0.67 $\pm$ 0.09 0.83 $\pm$ 0.04 0.00 $\pm$ 0.00
14 0.15 $\pm$ 0.03 0.13 $\pm$ 0.04 0.00 $\pm$ 0.00
15 0.68 $\pm$ 0.09 0.30 $\pm$ 0.08 0.00 $\pm$ 0.00
16 0.03 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
17 0.28 $\pm$ 0.08 0.02 $\pm$ 0.02 0.00 $\pm$ 0.00
18 0.10 $\pm$ 0.08 0.02 $\pm$ 0.02 0.00 $\pm$ 0.00
19 0.27 $\pm$ 0.16 0.58 $\pm$ 0.07 0.00 $\pm$ 0.00
Figure 10: Illustration of the selective weighting heuristic using (a) Embedding Distance Matrix (EDM) and (b) Embedding Distance Curve (EDC). In the demonstration, the robot successfully picks up a jar and places it into a basket. In the failed rollout, the robot fails during the picking stage, resulting in the absence of subsequent steps. The steps surrounding the picking procedure are identified as the Separation Segment and are assigned higher weights during adaptation to address the model’s shortcomings. Specifically, the Separation Segment is determined by the smoothed minimum $L_{2}$ distances from EDC—obtained from EDM, where each of its entry indicates the embedding distance between a demonstration and failed rollout frame, as shown in this figure.